OpenStudy (anonymous):

lim x=8 2-cubedroot x/ 8-x How to solve?

OpenStudy (anonymous):

\[\lim_{x\to 8}\frac{2-\sqrt[3]{x}}{8-x}\] ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i tried to solve it but it didn't work out

OpenStudy (anonymous):

ok and i assume you have to do this by hand, not l'hopital or recognizing it as a derivative right? then it will take some algebra to do this

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

if it was the square root you would multiply top and bottom by the conjugate, because \[a^2-b^2=(a+b)(a-b)\]you do something analagous here, using instead \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

in your case \(a=2,b=\sqrt[3]{x}\) so to get rid of the radical multiply top and bottom by \[4+2\sqrt[3]{x}+\sqrt[3]{x}^2\]

OpenStudy (anonymous):

you get \[\frac{(2-\sqrt[3]{x})(4+2\sqrt[3]{x}+\sqrt[3]{x^2})}{(8-x)(4+\sqrt[3]{x}+\sqrt[3]{x}^2)}\]

OpenStudy (anonymous):

wait, so i get the 4 because I square the 2, I get the value in the middle because i multiply the values in the first bracket together, and I get the third value by squaring the x cubed root.

OpenStudy (anonymous):

the numerator is \(8-x\) so you get \[\frac{1}{4+\sqrt[3]{x}+\sqrt[3]{x}^2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you are trying to get rid of the radical in the numerator you use \(a^3-b^3=(a-b)(a^2+ab+b^2)\) to do it, in other words \(a-b=2-\sqrt[3]{x}\)

OpenStudy (anonymous):

that way you will get rid of the cubed root up top

OpenStudy (anonymous):

ok, so i'm given the first bracket a-b

OpenStudy (anonymous):

right \(a-b=2-\sqrt[3]{x}\)

OpenStudy (anonymous):

oh, but should it not have the cubed root then? sorry, i know its a stupid question

OpenStudy (anonymous):

so in order to get \(8-x\) up top you multiply top and bottom by \((2^2+2\sqrt[3]{x}+\sqrt[3]{x}^2)\)

OpenStudy (anonymous):

you get \(2^3-\sqrt[3]{x}^3=8-x\) in the numerator

OpenStudy (anonymous):

which by some miracle cancels with the \(8-x\) in the denominator

OpenStudy (anonymous):

i am sure you ran across problem like this: \[\lim_{x\to 4}\frac{\sqrt{x}-2}{x-4}\] in that case it is enough to multiply top and bottom by \(\sqrt{x}+2\) because then the radical will be gone up top

OpenStudy (anonymous):

oh ok, i get how it works now. So if this was a square root i would multiply the top and bottom by (a+b) since it would be (a-b)

OpenStudy (anonymous):

that is using the fact that \((a-b)(a+b)=a^2-b^2\) you are doing the same thing here but you need to use the difference of two cubes, not the difference of two squares

OpenStudy (anonymous):

so you would get in my example \[\frac{\sqrt{x}-2)(\sqrt{x}+2)}{(x-4)(\sqrt{x+2})}\] \[=\frac{x-4}{(x-4)(\sqrt{x}+2)}\] \[=\frac{1}{\sqrt{x}+2}\]

OpenStudy (anonymous):

then replace \(x\) by \(4\) this situation is more complicated, but the idea is identical

OpenStudy (anonymous):

oh ok .

OpenStudy (anonymous):

second to last step is \[\frac{1}{4+\sqrt[3]{x}+\sqrt[3]{x}^2}\] last step is to replace \(x\) by \(8\)

OpenStudy (anonymous):

i think i get the idea now. So if it is a cubic i would multiply by (a^2 +ab+b^2) , it would either be positive or negative depending on (a-b). same idea if it was a square root, but i take the difference of squares.

OpenStudy (anonymous):

yes that is the entire idea

OpenStudy (anonymous):

ok. thank you so much

OpenStudy (anonymous):

by "cubic" you mean "cubed root" yw

OpenStudy (anonymous):

yea